OpenStudy (anonymous):

Which trinomial has a factor of (2x+3) A.2x^2+6x+9 B.2x^2+12x+9 C.4x^2+6x+9 D.4x^2+12x+9

OpenStudy (anonymous):

This may work try it. Multiply A x C Then find two numbers that can multiply into that but still give you b. For example in the first one multiply 2 x 9. then find two things that when multiplied equal 18, but add up to 6

OpenStudy (anonymous):

Is it 2x^2+12x+9 @giantsniners

OpenStudy (anonymous):

I need backup, I feel like my strategy does not work for this certain problem. @RadEn

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (helder_edwin):

if u put \(2x+3=0\) u get \(x=-3/2\). replace x=-3/2 in the polynomials, if it becomes zero, u have the answer

OpenStudy (anonymous):

okay thank you

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

I didnt get zero for any of them @helder_edwin

OpenStudy (shamil98):

4x^2+12x+9 = (2x+3)(2x+3) or (2x+3)^2

OpenStudy (anonymous):

So will it be 4x^2+12x+9 then? @shamil98

OpenStudy (shamil98):

yes.

OpenStudy (helder_edwin):

u should have gotten a zero when replacing x=-3/2 in (D). 4(-3/2)^2+12(-3/2)+9=4(9/4)-12(3/2)+9=9-18+9=0

OpenStudy (anonymous):

Okay thank you

OpenStudy (helder_edwin):

u r welcome